1. If f(x)=-x^2-4x+3, find f(-2)
2. Solve for x: 3(x-2)=-4(x+8)+10
3. State the domain of f(x)=x/x-4
4. Write the equation of the line whose slope is -3/7 and y-intercept is -4; write the answer in slope-intercept form.
text is College Algebra

Answer by edjones(7311)

f(x)=-x^2-4x+3
f(-2)=-4+8+3
=7
.
3(x-2)=-4(x+8)+10
3x-6=-4x-32+10
3x-6=-4x-22
7x=-16
x=-16/7
.
f(x)=x/x-4
Domain of x is all real numbers except 4. Division by zero not allowed.
.
m=-3/7
one point on the graph is (0,-4)
y-y[1]=m(x-x[1])
y-(-4)=-3/7(x-0)
y+8=-3x/7
y=(-3x/7)-8
.
Ed